### 16 Apr 2026 v2

# Visual Enhanced Depth Scaling for Multimodal Latent Reasoning

Abstract
Multimodal latent reasoning has emerged as a promising paradigm that replaces explicit Chain-of-Thought (CoT) decoding with implicit feature propagation, simultaneously enhancing representation informativeness and reducing inference latency. By analyzing token-level gradient dynamics during latent training, we reveal two critical observations: (1) visual tokens exhibit significantly higher and more volatile gradient norms than their textual counterparts due to inherent language bias, resulting in systematic visual under-optimization; and (2) semantically simple tokens converge rapidly, whereas complex tokens exhibit persistent gradient instability constrained by fixed architectural depths. To address these limitations, we propose a visual replay module and routing depth scaling to collaboratively enhance visual perception and refine complicated latents for deeper contextual reasoning. The former module leverages causal self-attention to estimate token saliency, reinforcing fine-grained grounding through spatially-coherent constraints. Complementarily, the latter mechanism adaptively allocates additional reasoning steps to complex tokens, enabling deeper contextual refinement. Guided by a curriculum strategy that progressively internalizes explicit CoT into compact latent representations, our framework achieves state-of-the-art performance across diverse benchmarks while delivering substantial inference speedups over explicit CoT baselines.
1 Introduction
(a) Token Frobenius Norm in Different Epochs
(b) Easy Samples
(c) Hard Samples
Figure 1: Panel (a) depicts the token-wise Frobenius norm of gradients within different layer throughout the overall training process. Notably, visual tokens (i.e., those with indices within about $\left[200,280\right]$) exhibit consistently larger gradient magnitudes accompanied by abrupt spikes, suggesting that they are more challenging to optimize compared to textual tokens (indices without $\left[200,280\right]$). Panels (b) and (c) further reveal distinct evolution patterns of the Nuclear norm in the QKVO projection matrix across layers and training epochs: gradients for easy samples decay rapidly and converge smoothly, whereas hard samples maintain elevated gradient norms with persistent oscillations. Here, $\mathbf{W}_{A}$ and $\mathbf{W}_{B}$ denote the low-rank matrix in LoRA [hu2022lora].
Over the past few years, large language models (LLMs) have achieved remarkable progress in complex reasoning, propelled by scaling laws in data volume and model capacity [DBLP:journals/corr/abs-2001-08361]. Advanced techniques such as Chain-of-Thought (CoT) prompting [zhang2023multimodal, shao2024visual] and reinforcement learning (RL) [deepseekr1] for trajectory optimization have proven highly effective in text-only domains. Extending these capabilities to the multimodal realm has thus become a pivotal research direction. Current approaches primarily follow three paradigms. First, Text-based Reasoning [zhang2023multimodal, MitraCCoT, lei2024scaffolding] generates explicit multi-step textual chains before producing an answer. However, these methods typically rely on static visual inputs, and recent studies [lookback2025, DBLP:journals/corr/abs-2502-17425] indicate that visual grounding deteriorates significantly over extended reasoning chains. Second, Tool-augmented Reasoning manipulates visual inputs through external operations (e.g., zooming or region enhancement) and injects intermediate visual hints into the reasoning trace. While powerful, these approaches are prone to redundant or invalid tool invocations, introducing noise that degrades performance and substantially increases inference latency. Recently, emerging researches position Latent Reasoning as a viable direction for optimizing multimodal reasoning. Unlike traditional methods that rely on explicit textual chains, latent reasoning encodes intermediate reasoning steps into compact continuous vectors. This paradigm offers compelling advantages, including higher inference efficiency, reduced annotation overhead, and the ability to learn dense, high-fidelity multimodal representations [zhang2023multimodal, DBLP:journals/corr/abs-2505-16552].
To gain deeper insights into the optimization dynamics of latent reasoning, we conducted a systematic analysis of gradient flows and parametric evolution during training. This investigation reveals two meaningful observations: (1) Visual-Text Optimization Disparity: Recent studies [DBLP:journals/corr/abs-2507-03019, DBLP:conf/iclr/0002KCI25] have highlighted the phenomenon of visual attention attenuation in MLLMs as the explicit Chain-of-Thought (CoT) reasoning chain extends. We discover that a similar degradation existing in latent reasoning. As depicted in Fig. 1(a), visual tokens consistently exhibit significantly higher gradient norms and pronounced volatility compared to textual tokens. We reckon that this disparity stems from a fundamental mismatch between continuous visual features and discrete text tokens. During joint training, textual modality dominates the optimization process, causing visual representations to undergo large and erratic updates as they struggle to establish the fine-grained visual-text alignment. (2) Fixed-Depth Optimization Dilemma: Beyond modality imbalance, we observe an architecture bottleneck in how models handle samples with different complexity. Partitioning the training data into easy (consistently correct) and hard (persistently incorrect) subsets based on early-stage validation accuracy, we tracked the gradient nuclear norms across the QKV and output projection matrices O (Fig. 1(b) and Fig. 1(c)). The trends reveal a stark divergence: easy samples exhibit smooth gradient decay, indicating stable convergence into favorable loss basins. In contrast, hard samples maintain persistently high gradient volatility even in late training epochs. This phenomenon underscores the necessity of iterative refinement for complex reasoning, which is crucial for parsing compositional patterns in complex contexts while mitigating inherent visual ambiguities in images. Fixed-depth architectures, however, lack the flexibility to adapt to varying token complexities, thereby trapping hard examples in oscillatory optimization trends.
Motivated by these observations, we propose a unified framework that strengthens fine-grained visual engagement and achieves token-wise depth scaling strategy to enable more precise and comprehensive contextual reasoning. Firstly, to mitigate visual optimization instability, we design a visual replay module, which dynamically replays the focused visual clues interleaved with thinking latents. This mechanism iteratively exposes and propagates key visual context across reasoning steps, fostering step-wise alignment with the target answer. Complementing this, we apply self-distillation supervision to enforce spatial coherence and preserve fine-grained visual details within the visual latents. Secondly, to address the fixed-depth optimization bottleneck, we propose a per-layer token router that dynamically allocates additional reasoning steps based on token complexity or information density. This design enables high-difficulty tokens to engage in prolonged contextual reasoning by reusing layer-wise knowledge, facilitating iterative representation refinement and adaptive prioritization of salient contextual cues. Finally, in contrast to methods [hao2024training, codi] that rely on knowledge distillation for direct latent supervision, we employ a curriculum learning strategy that progressively introduces latent tokens into the training pipeline. Extensive experiments show that our method can be effortlessly combined with various widely-used MLLM backbones to further enhance reasoning performance while maintain the satisfactory inference latency. The main contributions can be summarized as follows:
  • We systematically analyze token-level gradient dynamics during latent reasoning training, revealing two critical optimization bottlenecks: visual-text optimization disparity and fixed-depth optimization dilemma.
  • We present a unified curriculum-driven framework that progressively constructs interleaved latent representations. By integrating spatially-coherent visual constraints for fine-grained grounding and complexity-aware depth scaling, our approach enables robust and precise contextual reasoning.
  • Extensive experiments across twelve widely-used multimodal reasoning benchmarks demonstrate that our method achieves state-of-the-art performance while maintaining high inference efficiency.
2 Related Work
2.1 Explicit Multimodal Reasoning
Multimodal reasoning enables model to reason over information from different modalities to solve complex tasks. There are many prior works [hu2022promptcap, ssa, Zheng_NeurIPS2023, MitraCCoT, DBLP:journals/corr/abs-2601-01984, MitraCCoT, kamcot, dynfocus, super] focusing extensively on enhancing reasoning capabilities. Earlier work rely on the CoT prompting to perform explicit thinking steps in the text space before generating the final answer. However, this paradigm generally lack sufficient visual grounding capability and leads to unsatisfactory misalignment and hallucination [huang-etal-2024-visual, DBLP:journals/corr/abs-2404-18930]. To address these limitations, recent studies [Zheng_NeurIPS2023, hu2022promptcap, MitraCCoT] have explored converting visual information into textual formats prior to reasoning, leveraging external tools or specialized visual experts to generate descriptive representations that guide LLMs. For instance, hu2022promptcap pioneered the integration of visual captions, extracting semantic content as text and concatenating it with input prompts to bolster reasoning capabilities. To further enhance fine-grained reasoning, subsequent works have focused on regional understanding, aiming to improve textual expressiveness by describing specific image regions. Others [DBLP:journals/corr/abs-2601-01984, MitraCCoT, kamcot] identify entities and their relationships within images, facilitating fine-grained reasoning through explicit modeling of inter-entity connections.
A line of concurrent works advocates using vision-text interleaved format during the rationale generation and reasoning process. The model draws auxiliary lines or marks based on original image to record thinking path, zoom or crop regions, or perform code editing, etc. Building on these paradigms, zhang2023multimodal first proposed decoupling rationale generation from answer generation in the Vision-Text Reasoning field. Subsequently, Shao2024VisualCA annotates key regions of the original image in intermediate steps, training models to focus on image regions relevant to the answer. While some works [gao2024interleaved, zhang2025chain] further extract key image regions progressively during reasoning, combining visual information with textual reasoning to generate the final answer. Moreover, new methods [visualsketchpad, liu2025Visual] emulate human thought by sketching images during reasoning, focusing on core concepts, structures, and relationships while ignoring redundant details. Other works [li2025imaginereasoningspacemultimodal, chern2025thinking] generate new images with auxiliary markers during reasoning, combining them with text to improve reasoning in complex scenarios. To fully shift reasoning from the linguistic domain to the visual modality, xu2025visualplanningletsthink proposes to reasoning exclusively with dynamic generated images, demonstrating substantial performance gains in visual navigation tasks.
More recently, Vision-R1 [DBLP:journals/corr/abs-2503-06749] and VL-Rethinker [DBLP:journals/corr/abs-2504-08837] leverage Group Relative Policy Optimization [deepseekr1] (GRPO) to refine reasoning trajectories through rollout-based sampling and reward scoring. Complementing these policy-driven approaches, concurrent works further enhance reasoning capabilities via novel cognitive paradigms, including self-critiquing cycles [DBLP:journals/corr/abs-2405-06682, DBLP:conf/cvpr/CocchiMC0C25], iterative rethinking [lookback2025], and on-policy distillation [DBLP:journals/corr/abs-2601-18734].
2.2 Latent Reasoning
Different from explicit reasoning in the discrete token space, latent reasoning refers to internal computation performed in a hidden space before answer generation. hao2024training pioneers continuous latent space reasoning by feeding the last hidden states as input embeddings for the next step without generating intermediate discrete tokens, substantially reducing reasoning latency. However, subsequent studies have indicated that such paradigm may suffer from feature homogenization without explicit supervision on intermediate latent states. To address this limitation, a series of works attempt to enhance the quality of intermediate representations using diverse strategies. cheng2024compressed introduced variable-length contemplation tokens for latent reasoning, mitigating quality degradation caused by fixed-length constraints. Similarly, shen2025codi leveraged distillation tactic to align student and teacher hidden activations along with explicit supervision, thereby constraining latent reasoning paths. Beyond these efforts, DBLP:journals/corr/abs-2509-20317 adopted step-level supervision to further stabilize the reasoning space.
Recently, latent reasoning has been extended to Multimodal Large Language Models (MLLMs). Distinct from text-only LLMs, MLLMs necessitate the effective integration of visual features within the latent reasoning space. Several efforts [yang2025machine, li2025latent, pham2025multimodal, liu2025reasoningminddynamicmultimodal] have been dedicated to injecting visual cues into the latent space to facilitate visual-grounded reasoning. For instance, li2025latent emphasize image details by constructing a structured cognitive hierarchy, albeit relying on annotation-intensive multimodal reasoning data. Similarly, liu2025reasoningminddynamicmultimodal progressively select visual patches to inject into latent thinking tokens via confidence-guided policy gradient optimization. In this work, we systematically analyze gradient dynamics during latent reasoning training and reveal two critical bottlenecks towards token-wise optimization behavior.
3 Method
3.1 Preliminary: Implicit and Explicit Decoding
Given a multimodal input comprising a question $\mathcal{Q}$ and an image $\mathcal{V}$, we first tokenize them into a sequence of text embeddings $\mathbf{Q}=\{\mathbf{q}_{i}\}_{i=1}^{N_{q}}$ and visual features $\mathbf{V}=\{\mathbf{v}_{i}\}_{i=1}^{N_{v}}$ via a word embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{W}|\times d}$ and a pretrained visual encoder, respectively, where $|\mathcal{W}|$ denotes the vocabulary size and $d$ is the hidden dimension. Subsequently, we employ an autoregressive MLLM $\mathcal{F}_{\theta}$ to encode the concatenated input into an initial hidden state $\mathbf{H}^{(0)}\in\mathbb{R}^{P\times d}$, where $P=N_{q}+N_{v}$ represents the total number of tokens in the prefilling phase. During the decoding phase, we predetermine the number of implicit reasoning steps $T_{r}$ and explicit answer tokens $T_{a}$. The generation process consists of two following distinct stages.
Implicit Reasoning Phase. In the $t$-th implicit reasoning step, the model generates a continuous latent representation $\mathbf{z}_{t}$ conditioned on the original input sequence $\mathbf{X}=\left[\mathbf{Q}\,\|\,\mathbf{V}\right]$ and all preceding latent states. This iterative process is formulated as:
$\mathbf{z}^{(t)}=\mathcal{T}\left(\mathcal{F}_{\theta}\left(\mathbf{X}\,\|\,\mathbf{Z}_{<t}\right)\right),\quad t=1,\dots,T_{r},$ (1)
where $\mathbf{Z}_{<t}=\{\mathbf{z}^{(1)},\dots,\mathbf{z}^{(t-1)}\}$ denotes the sequence of latent tokens generated in previous steps, and $\mathcal{T}(\cdot)$ extracts the final vector representation from the model output (i.e., the hidden state corresponding to the last generated token). Unlike vanilla explicit reasoning, each step yields a informative continuous latent vector rather than a discrete token. The resulting latent sequence $\mathbf{Z}_{1:T_{r}}=\{\mathbf{z}^{(t)}\}_{t=1}^{T_{r}}$ is concatenated to the context before transitioning to the explicit answer decoding phase.
Explicit Answer Decoding. In the explicit phase, the model generates discrete answer tokens by sampling from the vocabulary distribution. The conditional probability of the $t$-th answer token $a_{t}$ is given by:
$p_{\theta}(a_{t}\mid\mathbf{X},\mathbf{Z}_{1:T_{r}},a_{<t})=\mathrm{Softmax}\left(\mathbf{W}_{o}\cdot\mathbf{h}_{t}^{\text{dec}}\right),$ (2)
where $\mathbf{h}_{t}^{\text{dec}}$ is the decoder hidden state at step $t$, and $\mathbf{W}_{o}\in\mathbb{R}^{|\mathcal{W}|\times d}$ is the output projection matrix that maps hidden states to vocabulary logits.
The likelihood of the complete answer sequence $\mathbf{a}_{1:T_{a}}$ is factorized as an autoregressive product:
$p_{\theta}(\mathbf{a}_{1:T_{a}}\mid\mathbf{X},\mathbf{Z}_{1:T_{r}})=\prod_{t=1}^{T_{a}}p_{\theta}(a_{t}\mid\mathbf{X},\mathbf{Z}_{1:T_{r}},a_{<t}).$ (3)
Here, the full context input for the decoder is defined as $\mathbf{U}_{t}=\left[\mathbf{X}\,\|\,\mathbf{Z}_{1:T_{r}}\,\|\,\mathbf{a}_{<t}\right]$, where $\|$ denotes sequence concatenation along the token dimension.
Figure 2: Left Panel: Schematic illustration of our framework. Input images are encoded into visual tokens by a pretrained visual encoder and projected into a text-centric semantic space aligned with the LLM, while questions are tokenized by the text tokenizer. Our method enhances a standard latent MLLM with two synergistic components during the implicit reasoning phase: (1) Spatially-Coherent Finer Visual Replay (SCF-VR): Attention-guided selection identifies salient visual regions at each implicit step, generating fine-grained visual latents enhanced by spatially-coherent constraints. (2) Routing Depth Scaling (RDS): A lightweight learnable router adaptively allocates additional reasoning steps for high-difficulty tokens or latents. Refined hidden states and visual latents are interleaved and propagated to subsequent reasoning steps. The overall objective is jointly optimized via standard cross-entropy loss and self-distillation loss. Right Panel: Detailed architecture of the SCF-VR module and token-wise depth scaling mechanism. TokenSR denotes the token super-resolution module, and Pool is the average pooling.
3.2 Spatially-Coherent Finer Visual Replay
As mentioned earlier, we empirically reveal the gradient disparities between visual and textual tokens throughout the learning dynamics. Specifically, visual tokens consistently exhibit substantially larger gradient norms and fluctuations compared to textual tokens, indicating that visual representations remain under-optimized despite their critical role in multimodal reasoning. Motivated by these insights, we introduce the visual replay module to reinforce the engagement of visual cues via salient region detection, while enhancing fine-grained spatially-coherent perception capabilities via self-distillation supervision at each reasoning step. Attention-Guided Region Focus. Several works [Li_2025_ICCV, DBLP:conf/naacl/ZhangQSYYXWGTY25] have demonstrated that LLMs exhibit fundamental visual grounding capabilities. To identify visually salient regions, we aggregate attention weights across all transformer layers and attention heads to obtain a consolidated spatial focus map. Specifically, given the $l$-th layer and the $h$-th attention head, we compute the mean attention map $\bar{\mathbf{A}}^{(t)}$ at reasoning step $t$:
$\bar{\mathbf{A}}^{(t)}=\frac{1}{L\cdot H}\sum_{l=1}^{L}\sum_{h=1}^{H}\mathbf{A}^{(l,h,t)},$ (4)
where $\mathbf{A}^{(l,h,t)}\in\mathbb{R}^{P^{(t)}\times P^{(t)}}$ denotes the attention matrix for layer $l$ and head $h$ at iteration $t$ ($1\leq t\leq T_{r}$), with $P^{(t)}$ representing the number of input tokens at iteration $t$. Here, $L$ and $H$ represent the total number of layers and heads, respectively. To obtain token-level attention scores, we extract the attention distribution from the most recently generated token to all preceding tokens via column-wise summation, i.e., $\mathbf{a}_{all}^{(t)}=\mathrm{colsum}(\bar{\mathbf{A}}^{(t)})\in\mathbb{R}^{P^{(t)}}$. Subsequently, we extract only the visual token attention scores from $\mathbf{a}_{all}^{(t)}$ using the image mask, denoted as $\mathbf{a}^{(t)}\in\mathbb{R}^{N_{v}}$, where $N_{v}$ is the number of visual tokens. This visual attention vector effectively captures which visual tokens are most relevant to the current reasoning context.
As visual focus evolves across reasoning steps, we iteratively select the top-$K$ attended visual tokens $\{\mathbf{v}_{i}^{(t)}\}_{i=1}^{K}$ as visual latents, which are integrated with hidden states $\mathbf{z}_{t}$. To prevent redundant re-selection and promote diverse exploration, we maintain a visited token set $\mathcal{V}^{(t)}_{\text{visited}}$, ensuring comprehensive visual coverage:
$\mathcal{I}^{(t)}=\operatorname{TopK}\left(\{a^{(t)}_{i}\mid i\in\mathcal{V}^{(t)}_{\text{visited}}\}\right)$, (5)
where $\mathcal{I}^{(t)}$ represents the indices of the $K$ visual tokens with the highest attention scores at step $t$, and $\mathcal{V}^{(t)}_{\text{visited}}$ is updated after each selection. The original embeddings of the selected tokens $\mathbf{V}_{\mathcal{I}^{(t)}}=\{\mathbf{v}_{i}\mid i\in\mathcal{I}^{(t)}\}$ are further weighted by normalized attention scores:
$\mathbf{B}^{(t)}=\mathrm{Diag}(\mathrm{Softmax}(\{a^{(t)}_{i}|i\in\mathcal{I}^{(t)}\}))\mathbf{V}_{\mathcal{I}^{(t)}},$ (6)
where $\mathrm{Diag}(\cdot)$ constructs a diagonal matrix from a vector.
Spatially-Coherent Regularization. Although leveraging learned attention within Transformers provides explainable visual locations, it often suffers from limited spatial continuity due to the scattered nature of selected tokens and introduces noise associated with the attention sink phenomenon [DBLP:conf/iclr/XiaoTCHL24]. To mitigate these issues and enhance fine-grained perception without external annotations, we introduce self-distillation supervision. This mechanism involves cropping the visual regions exhibiting spatial coherence, re-encoding them, and supervising the visual latents with these high-fidelity features.
Specifically, we first search for a $W\times W$ sub-grid patch that maximizes the density of attended visual tokens. Formally, we find the optimal top-left corner $(r^{*},c^{*})$ within the valid grid bounds:
$(r^{*},c^{*})=\underset{0\leq r,c\leq G-W}{\operatorname{argmax}}\;\mathcal{N}(r,c),$ (7)
where the density function $\mathcal{N}(r,c)$ counts the visited tokens falling within the window $\mathcal{N}(r,c)=\sum_{i\in\mathcal{I}^{(t)}}\mathbb{I}\big[r\leq r_{i}<r+W,c\leq c_{i}<c+W\big].$ Here, $(r_{i},c_{i})$ denotes the row-column position of token $i$ on the $G\times G$ grid, and $\mathbb{I}[\cdot]$ is the indicator function. This greedy selection ensures the cropped region captures a coherent visual context rather than scattered details. Then, the selected window is projected to pixel coordinates in the original image via mathematical transformation, cropped, and resized to the standard encoder input resolution using bilinear interpolation. We then re-encode this refined patch through the same visual encoder to obtain fine-grained representations $\{\mathbf{f}_{i}\}_{i=1}^{N_{v}^{\prime}}$. A global pooling operation $\mathrm{Pool}(\cdot)$ is applied to yield a robust reference token $\mathbf{u}^{\text{ref}}$:
$\mathbf{u}^{\text{ref}}=\mathrm{Pool}\big(\{\mathbf{f}_{i}\}_{i=1}^{N_{v}^{\prime}}\big).$ (8)
Finally, we align the coarse-grained global token $\mathbf{b}^{(t)}=\mathrm{Pool}(\mathbf{B}^{(t)})$ with this high-fidelity reference using a lightweight token super-resolution module $\mathcal{F}_{\text{SR}}:\mathbb{R}^{D}\to\mathbb{R}^{D}$. We minimize the reconstruction error as follows:
$\mathcal{L}_{\text{recon}}^{(t)}=\left\|\mathcal{F}_{\text{SR}}\left(\mathbf{b}^{(t)}\right)-\mathbf{u}^{\text{ref}}\right\|_{2}^{2}.$ (9)
This supervisory signal enables the model to prioritize spatially coherent and semantically intact visual contexts during latent generation through self-distillation.
3.3 Routing Depth Scaling
While the visual replay mechanism significantly enhances fine-grained grounding by introducing spatially-coherent visual latents, existing methods typically allocate uniform computational budgets across all tokens during contextual refinement. Our empirical analysis reveals a fixed-depth optimization dilemma, demonstrating that token representations exhibit heterogeneous optimization complexities during latent training, which necessitates adaptive reasoning depths. To address this limitation without modifying the pretrained VLM architecture, while effectively leveraging its inherent knowledge, we introduce a lightweight router that dynamically allocates additional reasoning steps exclusively to critical tokens at each iteration. Router Network. Let the input token sequence at the $t$-th reasoning step be denoted as $\mathbf{U}^{(t)}=[\mathbf{X}\|\mathbf{B}^{(1)}\|\mathbf{z}^{(1)}\|\cdots\|\mathbf{B}^{(t)}\|\mathbf{z}^{(t)}]\in\mathbb{R}^{P^{(t)}\times d}$, where $P^{(t)}$ represents the total sequence length and $d$ is the hidden dimension. To facilitate the illustration of our depth scaling mechanism, we decompose the forward pass into layer-wise computations. Specifically, within the $l$-th transformer layer, we first compute the intermediate feature $\mathbf{H}^{(l,t)}=f(\mathbf{U}^{(t)})\in\mathbb{R}^{P^{(t)}\times d}$, where $f(\cdot)$ denotes the transformation of the $l$-th transformer layer in LLM. Subsequently, a lightweight router network computes a scalar importance score for each token based on its corresponding hidden representation:
$\mathbf{s}^{(l,t)}=\mathcal{F}_{\text{router}}(\mathbf{H}^{(l,t)})\in\mathbb{R}^{P^{(t)}},$ (10)
[…]
Depth Scaling Computation. Obtaining the router weight, we select the top-$\alpha$ tokens for depth scaling, i.e., repeat iteration in one transformer block. For a given depth scaling step $d\in\{1,\ldots,D\}$, the token-wise representation update rule within a transformer layer can be generally formulated as follows:
$\mathbf{h}_{i}^{(d)}=\begin{cases}s^{(d)}_{i}\odot f_{\in T_{\alpha}(\mathbf{s})}(\mathbf{h}^{(d-1)}_{i},\mathbf{m}^{(d-1)}),&\text{if }i\in T_{\alpha}(\mathbf{s}^{(d)}),\\ \mathbf{h}_{i}^{(d-1)},&\text{if }i\notin T_{\alpha}(\mathbf{s}^{(d)}).\end{cases}$ (11)
For notational brevity, we omit the layer index $l$ and reasoning step $t$ in the following formulation, where $d$ denotes the refinement depth. For instance, the score vector $\mathbf{s}^{(l,t)}$ is simplified to $\mathbf{s}^{(d)}$. The function $f_{i\in T_{\alpha}(\mathbf{s}^{(d)})}(\cdot)$ indicates that the attention operation is restricted to the token subset $T_{\alpha}(\mathbf{s}^{(d)})$. Here, $\mathbf{h}^{(0)}_{i}$ represents the initial hidden state of the $i$-th token after the first forward pass, and $\mathbf{m}^{(d)}\in\mathbb{R}^{P^{(t)}\times P^{(t)}}$ denotes the attention mask corresponding to refinement step $d$. The condition $i\in T_{\alpha}(\mathbf{s}^{(d)})$ functions as a binary gating mechanism, ensuring that only tokens with importance scores exceeding the threshold undergo additional computation.
The router network dynamically determines whether to apply depth scaling based on the current contextual representations in a data-driven manner. Tokens identified as critical are iteratively processed through the transformation function for $d$ additional refinement steps, while non-critical tokens retain their prior representations to preserve computational efficiency. Finally, we aggregate the depth-wise refined representation with step-aware positional encoding to form the final hidden state incrementally:
$\mathbf{H}^{(D)}=\mathbf{H}^{(0)}+\sum_{d=1}^{D}\left(\mathbf{H}^{(d)}\odot\mathbf{e}^{(d)}\right).$ (12)
This adaptive scaling strategy effectively allocates deeper reasoning pathways to critical tokens, enabling the model to capture complex visual contexts and facilitate more profound reasoning capabilities.
Table 1: Comparison of various multimodal reasoning baselines across three benchmarks. We selected three datasets featuring detailed reasoning chains for training and evaluate on their test split, respectively. Three metrics are reported: Accuracy (%), Average number of Autoregressive Steps (# AR steps), and Average Generation Time (AVG. Time). Evaluations were conducted on the M3CoT, ScienceQA, and GQA benchmarks using Qwen2-VL-7B and Chameleon-7B. denotes the reimplementation for the methods with the same configuration with ours. No-CoT notes that directly predicts answers without generating intermediate steps.
M3CoT ScienceQA GQA
Method Model Acc.(%) $\uparrow$ # AR Steps $\downarrow$ Avg. Time $\downarrow$ Acc.(%) $\uparrow$ # AR Steps $\downarrow$ Avg. Time $\downarrow$ Acc.(%) $\uparrow$ # AR Steps $\downarrow$ Avg. Time $\downarrow$
No-CoT 45.4 64.4
Multimodal CoT [zhang2023multimodal] 42.5 106.3 3.10 58.3 83.9 2.44
CCoT [MitraCCoT] 44.1 177.2 5.31 63.8 164.0 5.23 51.2 76.4 7.21
ICoT [gao2024interleaved] 46.0 96.5 2.86 65.4 77.4 2.28
SCAFFOLD [lei2024scaffolding] 44.9 170.8 5.14 62.5 162.3 4.91 48.7 72.8 6.72
Chain-of-Focus [zhang2025chain] 64.3 185.7 2.63 91.2 162.3 2.09 61.8 128.6 3.01
IVT-LR 69.8 10.0 0.67 92.8 11.0 0.81 65.8 10.1 0.68
Ours Qwen2-VL 7B 73.0 7.0 0.86 95.9 7.2 1.02 67.4 9.2 0.82
No-CoT 28.4 48.5
Multimodal CoT [zhang2023multimodal] 30.6 110.5 3.62 50.7 98.7 3.33
CCoT [MitraCCoT] 31.4 168.4 5.35 51.3 174.2 5.39 33.1 150.6 5.31
ICoT [gao2024interleaved] 32.3 110.9 5.43 53.4 92.4 4.62
SCAFFOLD [lei2024scaffolding] 31.1 194.3 6.12 47.5 160.6 6.03 32.8 156.0 4.17
Chain-of-Focus [zhang2025chain] 36.5 739.4 3.09 61.2 717.1 2.56 34.6 360.4 2.98
IVT-LR 40.8 10.0 1.13 63.2 11.0 1.56 38.1 10.1 0.98
Ours Chameleon 7B 43.4 7.0 1.24 65.7 7.2 1.37 39.4 9.2 1.21
Model Data Size Vision-centric Reasoning Compositional Mathematics Reasoning
MMVP SeedBench-2-Plus HallusionBench HRBench BLINK MMStar MathVista MathVision $M$-Math
Zero-Shot VLMs
GPT-4o [gpt4v] 68.70 72.00 68.00 64.70 63.80 30.39 31.80
Qwen2.5-VL-7B [bai2025qwen2] 65.67 65.31 56.57 68.25 53.60 59.70 68.20 25.60 37.50
LLaVA-OneVision [li2024llava] 9M 74.00 61.22 51.10 63.00 49.34 59.13 58.60 - -
InternVL3.5-8B [wang2025internvl3] 70K 57.67 69.78 56.15 59.38 54.81 53.33 71.60 28.30 -
Explicit CoT Reasoning
Multimodal CoT [zhang2023multimodal] 68.10 54.11 63.60 - - 57.90 56.40 21.80 35.60
CCoT [MitraCCoT] 69.00 68.95 64.90 - - 58.70 57.80 22.50 36.30
ICoT [gao2024interleaved] 69.30 70.27 65.50 - - 60.40 58.90 23.30 37.00
Tool-use & RL Enhanced Reasoning
PAPO [wang2025perception] 39K 68.67 54.11 57.52 68.12 52.66 45.80 67.53 - -
Vision-R1 [huang2025vision] 200K 72.67 68.95 63.83 75.12 52.71 62.67 52.40 - 40.20
VL-Rethinker [wang2025vl] 39K 72.67 70.27 71.08 63.50 55.55 63.20 72.80 29.30 -
DeepEyes [zheng2025deepeyes] 47K 70.00 69.08 62.57 69.12 51.08 58.73 70.10 26.60 -
Latent Reasoning
LVR [li2025lvr] 470K 64.00 47.39 65.19 53.62 53.60 57.93 - - -
Monet [wang2025monet] 125K 68.00 65.88 56.36 68.00 50.71 60.33 - - -
DMLR [liu2025reasoningminddynamicmultimodal] 70.10 - 65.80 - 56.92 60.27 59.10 24.40 38.80
Laser [li2025latent] 267K 72.00 70.05 67.72 72.50 - 60.10 - - -
Ours 30K 76.67 66.86 68.63 74.21 57.96 60.82 69.80 25.89 39.82
Table 2: Performance comparison of our method against baselines across four paradigms: Zero-Shot VLMs, Explicit CoT, Tool-Use & RL, and Latent Reasoning. Benchmarks are categorized into three domains: Visual Perception (MMVP, Seed-Bench-2-Plus, HallusionBench, HR-Bench), Compositional Reasoning (BLINK, MMStar), and Mathematical Reasoning (MathVista, MathVision, $M$-Math). Among latent reasoning approaches, the best results are highlighted in bold, and the second-best are underlined. Our method is built upon the Qwen2.5-VL-7B backbone.
Table 3: We perform comprehensive ablation studies to analyze the contribution of each design choice, including hyper-parameter configurations and architectural components. Experiments are conducted on three CoT benchmarks with multiple backbone models. To save the training overhead, all models are trained for 16 epochs with a balanced sampled subset.
Method Model M3CoT $\uparrow$ ScienceQA $\uparrow$ GQA $\uparrow$
Base Qwen2-VL 2B 44.2 62.4 38.7
+RDS 47.6 (+3.4) 64.9 (+2.5) 40.4 (+1.7)
+RDS & SCF-VR 51.2 (+7.0) 66.3 (+3.9) 40.9 (+2.2)
Base Qwen2-VL 7B 63.7 78.2 52.3
+RDS 65.2 (+2.5) 80.5 (+2.3) 53.7 (+1.4)
+RDS & SCF-VR 66.4 (+3.6) 81.3 (+3.1) 54.1 (+1.8)
Base Qwen2.5-VL 3B 62.9 76.5 51.9
+RDS 64.5 (+1.6) 77.8 (+1.3) 52.9 (+1.0)
+RDS & SCF-VR 65.3 (+2.4) 78.5 (+2.0) 53.6 (+1.7)
Base Qwen2.5-VL 7B 71.1 85.7 56.3
+RDS 72.9 (+1.8) 87.2 (+1.5) 57.6 (+1.3)
+RDS & SCF-VR 73.9 (+2.8) 88.2 (+2.5) 58.3 (+2.0)
3.4 Training Procedure
Curriculum Latent Training. To mitigate annotation overhead and avoid the risk of human priors constraining model learning, we depart from prior approaches that rely on intermediate supervision for latent representations. Instead, we design a curriculum that facilitates the contextual and logical dependencies between implicit latents and explicit reasoning chains. In the initial stage, the model is trained with standard Chain-of-Thought (CoT) supervision, generating all reasoning steps explicitly to establish foundational reasoning capabilities. Subsequently, as training progresses, latent tokens are incrementally introduced. Specifically, one explicit reasoning step is progressively encapsulated into an informative $\left<\text{latent}\right$> token. Through this curriculum, each latent token progressively learns to ground the relevant contextual cues, effectively internalizing explicit reasoning chains into compact latent representations.
Training Objective. The overall training objective combines the standard language modeling loss with the self-distillation loss in the VR-SCF module,
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda\frac{1}{T_{r}}\sum_{t=1}^{T_{r}}\mathcal{L}_{\text{recon}}^{(t)},$ (13)
where $\mathcal{L}_{\text{CE}}$ is the standard cross-entropy loss over the language modeling task,
$\mathcal{L}_{\text{CE}}=-\sum_{t=1}^{T_{r}+T_{a}}\log p_{\theta}\left(a_{t}|\mathbf{U}^{(t)},a_{<t}\right),$ (14)
where $\lambda$ is a hyperparameter balancing the primary task and the auxiliary reconstruction objective. During inference, only the LLM component is active, and the visual play module is disabled, ensuring no additional computational overhead at test time.
4 Experiments
We first conducted extensive evaluations on twelve diverse benchmarks against state-of-the-art competitors to validate the superiority of our method. Then, we provide comprehensive analysis through multi-faceted ablation studies and in-depth investigations to elucidate the underlying mechanisms of our approach.
4.1 Experimental Setup
Evaluation Benchmark.
We evaluate our method on three tasks across twelve benchmarks: (1) Mathematics Reasoning (MathVista [mathvista], MathVision [mathvision], MM-Math [mmmath]); (2) Vision-centric Reasoning (Hallusion-Bench [hallusionbench], MMVP [mmvp], Seed-Bench-2-Plus [seedbench2plus], HR-Bench [hrbench], GQA [gqa]); (3) Multimodal Composition Reasoning (MMStar [mmstar], BLINK [blink], ScienceQA [lu2022learn], M3CoT [chen-etal-2024-m3cot]). Additional descriptive details can be found in the Appendix.
Baselines.
We evaluate our method against a comprehensive set of state-of-the-art baselines, which can be categorized into four paradigms: (1) Zero-shot VLMs (GPT-4o [gpt4v], LLaVA-OneVision [li2024llava], InternVL3.5-8B [wang2025internvl35], Qwen2.5-VL-7B [bai2025qwen25vl]), (2) Explicit CoT-based methods (SCAFFOLD [lei2024scaffolding], ICoT [gao2024interleaved], Multimodal-CoT [zhang2023multimodal], CCoT [MitraCCoT], Chain-of-Focus [zhang2025chain]), (3) Visual Enhanced methods, including tool-augmented reasoning (DeepEyes [zheng2025deepeyes]) and RL-enhanced VLM reasoning (Vision-R1 [huang2025vision], PAPO [wang2025perception], VL-Rethinker [wang2025vl]), and (4) Multimodal Latent Reasoning approaches (Laser [li2025latent], LVR [li2025lvr], Monet [wang2025monet], DMRL [liu2025reasoningminddynamicmultimodal]).
Training Dataset.
To facilitate reproducibility, we first detail the training data configuration. During the supervised fine-tuning (SFT) stage, we curate a subset of approximately 30K samples from OneThinker [DBLP:journals/corr/abs-2512-03043], selected for its diverse distribution of reasoning chain lengths. The statistical distributions of CoT length, reasoning steps, and topic categories within this subset are illustrated in Fig. 3.
Backbone Models.
To comprehensively evaluate the effectiveness and scalability of our approach, we instantiate our method across a diverse set of backbone architectures, including Qwen2-VL-2B/7B [wang2024qwen2], Qwen2.5-VL-3B/7B [bai2025qwen25vl], and Chameleon-7B [team2024chameleon].
Implementation Details.
All frameworks employ eager attention mode to enable explicit access to internal attention maps. We set the number of latent reasoning tokens to $T_{r}=4$, with $\alpha=32$ salient regions injected per refinement iteration. To balance computational efficiency and reasoning accuracy, we cap the maximum refinement depth at $D=1$. For the cosine annealing schedule, the region injection count decays from $\alpha_{s}=64$ to $\alpha_{e}=16$. Models are trained for 16 epochs by default.
We employ DeepSpeed ZeRO-2 optimization without CPU offloading, with a per-GPU batch size of 8. Training utilizes the Adam optimizer with a learning rate of $4\times 10^{-5}$ and $\beta_{1}=0.9$. For the proposed self-distillation loss, we set the weighting coefficient to $\lambda=1.0$ for 2B/3B-scale models and $\lambda=0.2$ for 7B-scale models, respectively. Input images are resized to different resolutions according to training dataset, please refer to Appendix for detailed preprocessing protocols. All experiments are conducted on 16 NVIDIA H20 GPUs (96GB VRAM each).
4.2 Overall Quantitative Results
Reasoning Accuracy.
As summarized in Table 1, we train and evaluate our method on three datasets that provide fine-grained CoT annotations. With the same training protocols, our approach achieves the best performance across both model backbones and all three datasets. Compared to traditional explicit CoT methods, our method yields clear-cut improvements, with an average accuracy gain of nearly 30% built upon the Qwen-2-VL architecture. Our approach further surpasses the second-best latent-based reasoning baseline (IVT-LR) by an average of +2.63% across all benchmarks.
As shown in Table 2, we further evaluate our method against widely-used mainstream benchmarks that diagnose multi-faceted reasoning capabilities. On most benchmarks, our approach achieves consistent improvements. Particularly on vision-centric benchmarks, our method attains remarkable gains in overall scores compared to previous state-of-the-art methods. Notably, we observe the most substantial improvement on MMVP (+4.67%), which employs CLIP-blind patterns and emphasizes the requirement for fine-grained visual discrimination. We attribute these gains to our designed RCF-VR. By maintaining a set of visual latents with fine-grained spatially-coherent constraints, our method effectively mitigates hallucination while capturing fine-grained visual details. Conversely, our approach exhibits modest performance on text-intensive benchmarks, notably SeedBench-2-Plus. This gap likely arises from the scarcity of domain-specific training data, underscoring the potential benefits of targeted fine-tuning for such tasks. Remarkably, our method achieves competitive performance against computationally intensive alternatives, including explicit Chain-of-Thought (CoT) and tool-augmented frameworks. Despite operating purely within the latent space—without relying on external knowledge retrieval or reinforcement learning-based optimization—our framework effectively captures rich semantic representations through token-wise depth scaling.
(a) CoT Length Distribution
(b) Thinking Step Distribution
(c) Topic Distribution
[…]
Reasoning Efficiency.
[…]
4.3 Ablation Analysis
Effect of Designed SCF-VR and RDS.
We introduce several model variants to validate the effectiveness of our proposed modules across four representative MLLM backbones. As summarized in Table 3, consistent performance improvements are observed across all architectures. Notably, when instantiated on Qwen2-VL-2B, our method achieves a substantial improvement of +7.0% on M3CoT. Furthermore, we observe that performance gains are less pronounced on the GQA benchmark. We hypothesize that this trend arises from the relatively straightforward nature of GQA, where both visual inputs and queries demand less complex reasoning. Consequently, baseline models can adequately address the task requirements, rendering the contributions of our designed modules less impactful in such scenarios.
Table 4: Performance comparison with different mapping networks of $\mathcal{F}_{\text{SR}}$ on four kinds of backbones.
M3CoT (Acc %)
Method Linear Proj. Vanilla Conv. Point-wise Conv. [DBLP:journals/corr/abs-1712-05245]
Qwen2-VL-2B 51.2 43.2 43.7
Qwen2-VL-7B 66.4 61.4 62.3
Qwen2.5-VL-3B 65.3 61.1 62.6
Qwen2.5-VL-7B 73.9 69.2 69.8
Dataset M3CoT ScienceQA GQA
(a) Curriculum Training
Base 62.0 75.4 51.3
Base + Curr. 62.9 (+0.9) 76.5 (+1.1) 51.9 (+0.6)
(b) Token Selection in Router $T_{\alpha}(\mathbf{s})$
$\alpha=32$ 62.9 76.5 51.9
$\alpha=16$ 62.1 (-0.8) 75.3 (-1.2) 51.2 (-0.7)
$\alpha=64$ 62.8 (-0.1) 76.2 (-0.3) 51.9 (-0.0)
Cosine-annealed Retention 62.6 (-0.3) 76.5 (-0.0) 51.7 (-0.2)
(c) Selection Strategy
32-Patch 62.9 76.5 51.9
16-Patch 61.7 (-1.2) 75.5 (-1.0) 51.3 (-0.6)
Soft-Mix 62.2 (-0.7) 75.8 (-0.7) 51.9 (-0.0)
(d) Position Encoding
Keep Old Position 62.9 76.5 51.9
Rearrange Position 62.2 (-0.7) 76.5 (-0.0) 51.3 (-0.6)
Table 5: Ablation Experiments. We provide ablation analysis of key parameters and experimental settings on three benchmarks: M3CoT, ScienceQA, and GQA. All the variants adopt Qwen2.5-VL-3B as the base model. Base refers the model that adopts RDS module and SCF-VR module.
Effect of Curriculum Training.
Curriculum training is critical to our method design. As evidenced in Table 5(a), incorporating the progressive training strategy yields a 0.9% performance improvement over the baseline. This paradigm facilitates the gradual integration of latent representations into the optimization process, enabling each latent to progressively ground relevant contextual cues.
Effect of Retention Parameter $\alpha$.
For the parameter $\alpha$ in $T_{\alpha}(\cdot)$, we compare different retention strategies. As illustrated in Table 5, we evaluate fixed top-$\alpha$ settings with $\alpha\in\{16,32,64\}$ and a cosine-annealed token retention schedule (CTR), which gradually reduces the token retention ratio layer by layer following a cosine schedule. Formally, for a model with $L$ layers, the retention ratio for the $l$-th layer is,
$\alpha(l)=\mathrm{round}\left(\alpha_{s}+\frac{\alpha_{s}-\alpha_{e}}{2}\left[1+\cos\left(\frac{\pi l}{L}\right)\right]\right),$ (15)
where $\alpha_{s}$ and $\alpha_{e}$ are two endpoints that enables flexible control over computational cost. As can be seen panel (b) in Table 5, results show that fixed top-$\alpha$ with $\alpha=32$ achieves optimal performance, while cosine-annealed schedules yield consistently inferior results. We attribute this result to that: as visual tokens and latent hidden states are gradually appended to the token sequence, larger contextual scopes becomes essential for deeper token interaction. Consequently, strategies that progressively decrease the token retention ratio during depth scaling unavoidably limit these critical interactions, thereby hindering the model’s capacity to develop comprehensive contextual understanding.
Visual Latents Formation.
The Soft-Mix strategy aggregates 32 selected patches into a single visual latent token via weighted summation, guided by reweighted attention scores. In contrast, the 32-patch and 16-patch strategies directly utilize the top 32 and 16 patches with the highest attention scores as visual latents at each reasoning step, respectively. As summarized in Table 5, the 32-patch strategy yields superior performance. Notably, Soft-Mix achieves performance comparable to 16-patch while utilizing only a single compact representation. We attribute this performance to the expanded search scope provided by the 32-patch strategy, which offers the router a richer pool of visual tokens, thereby effectively raising the upper bound of depth scaling capabilities.
4.4 In-depth Analysis
Impact of $\lambda$.
As shown in Figure 4, model performance exhibits high sensitivity to variations in the hyperparameter $\lambda$. Notably, larger models (e.g., 7B) achieve peak performance at $\lambda=1.0$, whereas the 2B model attains optimal results at a substantially lower value of $\lambda=0.2$. We attribute this divergence to distinct capacity across model scales: smaller models are primarily bottlenecked by visual perception capabilities, thus requiring larger degree of visual play to establish effective visual grounding for downstream reasoning. In contrast, larger models possess sufficient grounding capability, and they benefit from a more balanced allocation between visual and linguistic signals, avoiding over-reliance on visual features that could otherwise suppress the development of complex reasoning chains.
Figure 4: Sensitivity analysis of $\lambda$.
Figure 5: (a) and (b) illustrate the performance with different number of event prototypes and different ratio of filtered event prototypes, respectively.
Impact of different window size $W$.
As shown in Table 6, $W$ = 3 achieves optimal performance. We reckon that smaller window sizes may fail to comprehensively cover salient visual regions, whereas larger window sizes, while providing broader contextual information, tend to introduce excessive visual noise.
Figure 6: Visualization of cropped region in each latent reasoning step. Dotted line denotes the correspondence between reasoning rationales and cropped region.
Latent Behavior Analysis.
As depicted in Figure 5, we visualize the embedding space distribution of multimodal features for both the baseline and our method. The latent tokens learned by our approach form several distinct clusters, situated centrally within the text embedding manifold. Notably, compared to the baseline, our latents exhibit closer proximity to visual embeddings. This observation suggests that our latent tokens encapsulate richer reasoning semantics while facilitating deeper integration of visual information.
Table 6: Performance comparison of different sub-grid window sizes on the M3CoT benchmark. We set $\lambda=1.0$ for the 2B/3B models and $\lambda=0.2$ for the 7B model. Here, $W$ denotes the sub-grid window size, with $W=10$ is the window size of extracted visual feature.
M3CoT (Acc.%)
Method Grid Size $W$ = 2 $W$ = 3 $W$ = 5
Qwen2-VL-2B 10 $\times$ 10 50.8 51.2 50.3
Qwen2-VL-7B 65.7 66.4 65.2
Qwen2.5-VL-3B 65.0 65.3 64.5
Qwen2.5-VL-7B 73.4 73.9 72.1
Visualization of Crop Region.
As illustrated in Figure 6, we visualize the attention-guided crops across successive reasoning steps. Taking the third case as a representative example, which queries the purpose of the train cart, the model initially localizes the region corresponding to the description “shape like a house” during the first reasoning step. In subsequent steps, the attention focus progressively shifts toward the “entrance door”. Remarkably, without any explicit fine-grained supervision, our curriculum training paradigm enables latent tokens to progressively capture logical dependencies and establish robust visual-semantic alignments. This empirically demonstrates the model’s capacity to jointly perform multi-step visual grounding and generate coherent reasoning chains.
Impact of Position Encoding.
Regarding positional encoding during depth scaling, we evaluate two strategies. The Keep Old Position variant preserves the original relative positional embeddings of the selected tokens. In contrast, the Rearrange Position variant sequentially re-indexes the selected tokens from $1$ to $K$ according to the reduced sequence length. As demonstrated in Table 3, preserving the original positions yields superior performance. Although re-indexing may enhance local contextual modeling among salient tokens, it risks introducing positional inconsistencies with embeddings from preceding reasoning steps and discards inherent structural priors, ultimately leading to suboptimal results.
5 Conclusion
In this paper, we empirically reveal two critical observations: the vision-text optimization disparity and the fixed-depth optimization dilemma. In light of these findings, we propose a visual replay module and routing depth scaling to collaboratively enhance visual perception and exploit critical tokens for deeper contextual reasoning. Instead of relying on extensive latent-supervised annotations or predefined visual priors, we adopt curriculum training to progressively capture rich contextual information, yielding highly informative latent representations. Extensive experiments demonstrate that: (1) our approach not only reduces decoding latency but also achieves superior performance across diverse benchmarks; (2) the proposed components exhibit strong generalization across various backbones and tasks; and (3) visualization analysis reveals significantly enhanced visual grounding of the learned latent tokens. In future work, we plan to extend this framework to more complex reasoning scenarios, such as long-term video understanding, and scale it to larger model architectures.